Q2. (c) Find the general solution of the first-order linear partial differential equation xux + yuy = u and investigate first-order, quasi-linear and linear partial differential equations. 



1
Expert's answer
2021-02-24T07:39:14-0500

We have the linear differential equation

"xu_x+yu_y=u"


The system of eqautions is,

"\dfrac{dx}{x}=\dfrac{dy}{y}=\dfrac{du}{u}"


Now, taking first two terms

"\dfrac{dx}{x}=\dfrac{dy}{y}"


Integrating Both the sides-


"\text{lnx}=\text{lny}+C_1\\

\text{lnx}-\text{lny}=\text{ln}e^{C1}"


"(\dfrac{x}{y})=e^{C1}\\

\dfrac{x}{y}=C1'\\

\phi=\dfrac{x}{y}=C1'"


Now, taking last two terms


"\Rightarrow\dfrac{dy}{y}=\dfrac{du}{u}\\\Rightarrow

lny=lnu+C_2\\

\Rightarrow lny=lnu+lne^{C_2}\\\\

\Rightarrow ln(\dfrac{y}{u})=lne^{C_2}"


"\Rightarrow \dfrac{y}{u}=e^{C2}"


"\Rightarrow \dfrac{y}{u}=C_2'"


"\Rightarrow\epsilon=\dfrac{y}{u}=C_2'"


Hence, the general solution is "f(\phi,\epsilon)=f(\dfrac{x}{y},\dfrac{y}{u})=0"